OpenStudy (anonymous):

Together, two sports clubs (A and B) have x members; A has a members and B has b members. Some of the persons are members of both sports clubs. Which of the following expressions describes how many persons are members in only one of the two sports clubs? Together, two sports clubs (A and B) have x members; A has a members and B has b members. Some of the persons are members of both sports clubs. Which of the following expressions describes how many persons are members in only one of the two sports clubs? @Mathematics

OpenStudy (anonymous):

x=a+b-k where k denote number of persons joined both clubs then x-k is number of persons join only a club x-k= a+b-2k

OpenStudy (anonymous):

Thanks